# Exploring Pessimism and Optimism Dynamics in Deep Reinforcement Learning

### v1

@@ 4 Utility Soft Actor-Critic (USAC): framework and algorithms
where $U^{\mathcal{Q}}_{\lambda}(s,a)=\mathbb{E}_{Q\sim\mathcal{Q}}[Q(s,a)]$ when $\lambda=0$. This utility function is commonly used for measuring risk in finance, economics, and decision-making under uncertainty [9, 34]. [^3] The utility function in 8 is also referred to as expected utility, exponential utility, exponential risk measure, generalized mean, or entropic risk measure according to the context [8, 9]. It also shares similarities with the utility functions employed in risk-sensitive RL [14, 26, 22] and distributional RL [24, 1, 2]. It’s important to note that we apply this utility function to the $\mathcal{Q}$ distribution rather than the return distribution, as is typically done in risk-sensitive and distributional RL.
@@
In this work, we use 8 as a measure of pessimism or optimism of the $\mathcal{Q}$ distribution: a positive value of $\lambda$ implies an optimistic view, while a negative value indicates a pessimistic view. This interpretation is clear from the Taylor expansion: $U^{\mathcal{Q}}_{\lambda}=\mathbb{E}_{Q\sim\mathcal{Q}}[Q]+\frac{\lambda}{2}\mathbb{V}_{Q\sim\mathcal{Q}}[Q]+\mathcal{O}(\lambda^{2})$. Consequently, the utility function $U^{\mathcal{Q}}_{\lambda}$ spans the entire spectrum of the $\mathcal{Q}$ distribution by varying $\lambda$. Another important property of the utility is that when $\mathcal{Q}$ is a Dirac delta distribution centered at $Q^{\pi}$, the $Q$-function of a given policy $\pi\in\Pi$, the expected utility function $U^{\mathcal{Q}}_{\lambda}$ is equivalent to $Q^{\pi}$ for any finite $\lambda$.
4.1 USAC framework: policy evaluation, improvement and guarantees
In this section, we will explain the policy evaluation and policy improvement steps, along with the corresponding guarantees for our USAC framework.
[…]
@@ 4.1 USAC framework: policy evaluation, improvement and guarantees
$T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a)=r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[U_{\lambda}^{\mathcal{Q}}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime}))\right]-U_{\lambda}^{\mathcal{Q}}(s,a)$
should be diminished. Nevertheless, it fosters exploration during learning by maintaining a distribution over the $Q$-functions, which can guide exploration.
@@
On the other hand, given a $\mathcal{Q}$ distribution associated with the current policy $\pi$, denoted as $\mathcal{Q}^{\pi}$, we update the policy, similar to 4, through the following maximization problem:
@@
4.2 USAC algorithm
The USAC learning process follows the general actor-critic algorithm. In particular, we estimate the desired quantities by samples from a replay buffer $(s,a,r,s^{\prime})\sim\mathcal{D}$. As the operator $T^{\pi}$ should bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed-point $Q^{\pi}$, the USAC framework learns the critic $Q$ (an estimate of $Q^{\pi}$) as in 2 but with the critic’s target value defined by
$y(s,a,r,s^{\prime})=r+\gamma[U_{\lambda_{\text{critic}}}^{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}),$ (9)
which is built using a target critic distribution $\bar{\mathcal{Q}}$ with some utility parameter $\lambda_{\text{critic}}$ chosen specifically for the critic. Similar to 4, the actor learns its policy by
@@ 4.2 USAC algorithm
Note that although Moskovitz et al. [20] consider an optimism/pessimism parameter that can switch between two values over time, they use the same parameter for both actor and critic training. In contrast, our work decouples the optimism/pessimism levels of the actor and critic. This novel approach allows us to select a pessimistic parameter $\lambda_{\text{critic}}$ for learning the critic while using an optimistic parameter $\lambda_{\text{actor}}$ for training the actor. This strategy effectively addresses the overestimation bias observed during critic training and mitigates the issue of pessimistic under-exploration encountered by the actor. Consequently, our approach promotes a stable critic and an explorative actor simultaneously.
Later in the experiments presented in Section 5, we will explore various combinations of pessimism and optimism levels. Before proceeding, we will ground the algorithm by considering a specific distribution. It’s worth noting that while we focus on the Laplace distribution in this paper, the framework described in this section is applicable to other distributions as well. The specific choice of Laplace is simply because it makes it more tangible to fine-tune the level of pessimism/optimism in the critic and actor, and to compare with other actor-critic algorithms.
4.3 USAC algorithm: Laplace distribution
The utility formulation in 8 enables us to compute the utility for any distribution with a moment-generating function. Let’s consider a specific case where $\mathcal{Q}$ follows a Laplace distribution. The algorithm overview is provided in Algorithm 1. Here, we denote the utility parameter associated with this distribution as $\kappa$, reflecting its characteristics within the Laplace context.
@@ 4.3 USAC algorithm: Laplace distribution
$g(\kappa)=\begin{cases}\log(1/(1-\kappa^{2}))/\sqrt{2}\kappa&\>\text{for}\>\kappa\in(-1,1)\backslash\{0\},\\ 0&\>\text{for}\>\kappa=0.\end{cases}$ (12)
The proof of Proposition 1 can be found in Appendix A. Additionally, we present a corresponding proposition for Gaussian-distributed $Q$-values in Appendix B. In Proposition 1, we observe that by varying $\kappa$ in the small interval $(-1,1)$, $U^{\mathcal{Q}}_{\kappa}$ spans the entire spectrum of the $\mathcal{Q}$ distribution. Specifically, as $\lim_{\kappa\rightarrow-1}g(\kappa)\rightarrow-\infty$, $\lim_{\kappa\rightarrow 0}g(\kappa)\rightarrow 0$, and $\lim_{\kappa\rightarrow 1}g(\kappa)\rightarrow\infty$.
Now suppose the uncertainty follows a Laplace distribution. Following the algorithmic steps outlined in Section 4.2, we can estimate the utility of the target critic distribution $\bar{\mathcal{Q}}$ by two target critics $\{\bar{Q}_{k}\}_{k=1,2}$, which are delayed versions of the critics $\{Q_{k}\}_{k=1,2}$. The utility function is given by $U^{\bar{\mathcal{Q}}}_{\kappa_{\text{critic}}}=\mu_{\bar{\mathcal{Q}}}+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}$ with utility parameter $\kappa_{\text{critic}}\in(-1,1)$, where
$\mu_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}(\bar{Q}_{1}(s,a)+\bar{Q}_{2}(s,a))\quad\text{and}\quad\sigma_{\bar{\mathcal{Q}}}(s,a)=\frac{1}{2}\lvert\bar{Q}_{1}(s,a)-\bar{Q}_{2}(s,a)\rvert.$ (13)
Hence, the target value to train the critic becomes
$y(s,a,r,s^{\prime})=r+\gamma[\mu_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],\quad a^{\prime}\sim\pi(\cdot|s^{\prime}).$ (14)
Similarly, the actor learns its policy by
$\operatorname*{arg\,max}_{\pi\in\Pi}\mathbb{E}_{a\sim\pi(\cdot|s)}[\mu_{\tilde{\mathcal{Q}}}(s,a)+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s,a)-\alpha\log\pi(a|s)],\quad\kappa_{\text{actor}}\in(-1,1),$ (15)
where $\mu_{\tilde{\mathcal{Q}}}$ and $\sigma_{\tilde{\mathcal{Q}}}$ are estimated in similar manner as 13 on $\tilde{\mathcal{Q}}$ using the critics $\{Q_{k}\}_{k=1,2}$. The pseudo-code of this algorithm can be found in Algorithm 1.
We note that a smaller $\kappa$ value indicates more pessimism, while a larger $\kappa$ value indicates more optimism. By choosing $\kappa$ such that $g(\kappa)=-1$ (i.e., $\kappa\approx-0.83$), the utility in 11 simply reduces to the minimum of the two critics. Thus, choosing $\kappa_{\text{critic}}=\kappa_{\text{actor}}\approx-0.83$, we obtain the pessimism choices of SAC [12] and TD3 [10]. On the other hand, by choosing $\kappa_{\text{critic}}\approx-0.83$ and a positive $\kappa_{\text{actor}}$, we have a pessimism critic and an optimistic actor as explored in OAC [7]. Lastly, by letting a multi-armed bandit algorithm choose $\kappa_{\text{critic}}=\kappa_{\text{actor}}\in\{-0.916563,0\}$, we recover the pessimistic/optimistic design suggested by TOP [20]. [^4] To recover the pessimistic case of the TOP algorithm, we can choose $\kappa=-0.916563$, which corresponds to having the function $g$ in 12 equal to $\sqrt{2}$. We note that by choosing $-0.916563$, TOP considers a more pessimistic critic target than SAC, TD3, and OAC.
In Section 5, we will explore various $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ pairs and discuss whether the previous wisdom about the effectiveness of these strategies holds across different environments.
Algorithm 1 USAC with Laplace distribution (Section 4.3)
1:Input: Averaging parameter $\tau\in(0,1)$, learning rates $\eta_{\theta},\eta_{\phi}>0$, mini-batch size $n\in\mathbb{N}$, pessimism/optimism coefficients $\kappa_{\text{actor}},\kappa_{\text{critic}}\in(-1,1)$
2:Initialize: Replay buffer $\mathcal{D}=\varnothing$, initial state $s_{0}\sim p_{0}$, critic $\{\theta_{k}\}_{k=1,2}$, target critic $\{\bar{\theta}_{k}\}_{k=1,2}$ and actor policy $\phi$
3:for each time step do
4:     for each environment step do
5:         $a_{t}\sim\pi_{\phi}(\cdot|s_{t})$triangleright Sample action from the policy $\phi$
6:         $s_{t+1}\sim p(\cdot|s_{t},a_{t})$triangleright Sample transition from the environment
7:         $\mathcal{D}\leftarrow\mathcal{D}\cup(s_{t},a_{t},r_{t},s_{t+1})$triangleright Store transition in the replay pool
8:     end for
9:     for each training step do
10:         $\{(s_{i},a_{i},r_{i},s_{i}^{\prime},a_{i}^{\prime}):(s_{i},a_{i},r_{i},s_{i}^{\prime})\sim\mathcal{D},a_{i}^{\prime}\sim\pi_{\phi}(\cdot|s_{i}^{\prime})\}_{i=1}^{n}$triangleright Sample mini-batch
11:         for $i\in\{1,\dots,n\}$ do $\triangleright$ Compute target critic distribution $\bar{\mathcal{Q}}$
12:              $\mu_{\bar{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}(Q_{\bar{\theta}_{1}}(s_{i},a_{i})+Q_{\bar{\theta}_{2}}(s_{i},a_{i}))$
13:              $\sigma_{\bar{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}\lvert Q_{\bar{\theta}_{1}}(s_{i},a_{i})-Q_{\bar{\theta}_{2}}(s_{i},a_{i})\rvert$
14:         end for
15:         for each critic, $k\in\{1,2\}$ do
16:              $\theta_{k}\leftarrow\theta_{k}-\eta_{\theta}\nabla_{\theta_{k}}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(Q_{\theta_{k}}(s_{i},a_{i})-y(s_{i},a_{i},r_{i},s_{i}^{\prime})\right)^{2}\right\}$triangleright Update critic
17:             with $y(s_{i},a_{i},r_{i},s_{i}^{\prime})=r_{i}+\gamma[\mu_{\bar{\mathcal{Q}}}(s_{i}^{\prime},a_{i}^{\prime})+g(\kappa_{\text{critic}})\sigma_{\bar{\mathcal{Q}}}(s_{i}^{\prime},a_{i}^{\prime})-\alpha\log\pi_{\phi}(a_{i}^{\prime}|s_{i}^{\prime})]$
18:         end for
19:         for $i\in\{1,\dots,n\}$ do $\triangleright$ Compute actor distribution $\tilde{\mathcal{Q}}$
20:              $\mu_{\tilde{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}(Q_{\theta_{1}}(s_{i},a_{i})+Q_{\theta_{2}}(s_{i},a_{i}))$
21:              $\sigma_{\tilde{\mathcal{Q}}}(s_{i},a_{i})\leftarrow\frac{1}{2}\lvert Q_{\theta_{1}}(s_{i},a_{i})-Q_{\theta_{2}}(s_{i},a_{i})\rvert$
22:         end $\triangleright$ Update actor
23:         $\phi\leftarrow\phi+\eta_{\phi}\nabla_{\phi}\left\{\frac{1}{n}\sum_{i=1}^{n}\left(\mu_{\tilde{\mathcal{Q}}}(s_{i},a_{i})+g(\kappa_{\text{actor}})\sigma_{\tilde{\mathcal{Q}}}(s_{i},a_{i})-\alpha\log\pi_{\phi}(a_{i}|s_{i})\right)\right\}$
24:         for each critic, $k\in\{1,2\}$ do
25:              $\bar{\theta}_{k}\leftarrow\tau\theta_{k}+(1-\tau)\bar{\theta}_{k}$triangleright Update target critic
26:         end for
27:     end for
28:end for
5 Experiments
The goal of our experiments is to explore the interplay between critic and actor pessimism/optimism, estimation error (i.e., overestimation/underestimation), and performance. We compare our USAC algorithm with Laplace distribution (Section 4.3) with prior off-policy actor-critic algorithms. Specifically, we consider SAC [12], TD3 [10], OAC [7] and TOP [20] as baselines, as they represent the most relevant algorithms that incorporate optimism and pessimism. For SAC and USAC, we use auto-tuned entropy temperature $\alpha$ [13], while for OAC and TOP, we adopt the best parameters reported in their respective papers.
We conduct our experiments on 5 continuous control environments in the MuJoCo physics engine [30]. Each experiment is repeated across 5 different seeds, each comprising one million time steps, and evaluated over 10 episodes. The reported results represent the average (along with the standard deviation) across these repetitions. Additional details about architectures, baselines, computational cost, environment properties, grid search, hyper-parameters, learning curves, area under learning curves, estimation error plots and their values can be found in Appendix C.
Insights on USAC performance
In Table 1, we present two sets of results for our USAC algorithm: a default and a best pair of parameters $(\kappa_{\text{critic}},\kappa_{\text{actor}})$; we will discuss how these parameters are selected below. These results demonstrate that our USAC algorithm can outperform (or match) all baseline algorithms across all environments. This success reaffirms our hypothesis, as outlined in Section 4, that the levels of pessimism and optimism should vary between the critic and actor, as well as across environments [20]. In particular, our findings reveals that there exist pairs of utility parameters, $(\kappa_{\text{critic}},\kappa_{\text{actor}})$, capable of overcoming overestimation while facilitating exploration, thus enhancing overall performance. In addition, USAC smaller standard deviation indicates an improved model stability during learning. The corresponding learning curves and estimation error curves for these results are presented in Figures 5 and 6 (Appendix C).For the Humanoid-v4 environment, we found that the default parameter pair outperformed the best pair identified through grid search, e.g., see Figure 2. Therefore, we used the default pair for this environment in our results (Table 1). However, we conjecture that by expanding our (limited) grid search, a better pair may exist.
Table 1: Final return on MuJoCo environments trained with 1M time steps, averaged over 5 seeds. The best algorithms are highlighted in bold. $\pm$ corresponds to the standard deviation across repetitions. The default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ are both $-0.831559$. The best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ are listed in Table 2.
Environment USAC (ours) SAC TD3 OAC TOP
Default $(\kappa_{\text{critic}},\kappa_{\text{actor}})$ Best $(\kappa_{\text{critic}},\kappa_{\text{actor}})$
Ant-v4 $\boldsymbol{5139\pm 978}$ $\boldsymbol{5158\pm 1186}$ $4756\pm 1411$ $4091\pm 1303$ $4177\pm 1392$ $4334\pm 1276$
HalfCheetah-v4 $\boldsymbol{11024\pm 849}$ $\boldsymbol{11736\pm 317}$ $10763\pm 895$ $10570\pm 1400$ $8684\pm 1678$ $7311\pm 3074$
Hopper-v4 $3194\pm 810$ $\boldsymbol{3442\pm 247}$ $3185\pm 537$ $1986\pm 1154$ $3293\pm 113$ $\boldsymbol{3367\pm 154}$
Humanoid-v4 $\boldsymbol{5602\pm 505}$ $5602\pm 505^{\star}$ $\boldsymbol{5503\pm 373}$ $5149\pm 737$ $5390\pm 234$ $5332\pm 445$
Walker2d-v4 $\boldsymbol{4525\pm 534}$ $\boldsymbol{4530\pm 767}$ $3757\pm 1282$ $4369\pm 601$ $3467\pm 1200$ $4317\pm 631$
Exploring the dynamics of pessimistic and optimistic with USAC
To explore the dynamics of pessimism and optimism in USAC, we conducted a grid search across the 5 continuous control environments in MuJoCo. More details about this grid search can be found in Appendix C. We examined critics ranging from pessimistic to semi-optimistic and actors from pessimistic to optimistic. Specifically, we used $\kappa_{\text{critic}}\in\{-0.831559,-0.5,-0.33\}$ [^5] Recall that $\kappa=-0.831559$ corresponds to $g(\kappa)=1$, simplifying the utility function in 11 to the minimum of the two critics. Therefore, by setting $\kappa_{\text{critic}}\approx-0.83$, we align with the pessimistic choices of SAC [12], TD3 [10], and OAC [7]., and $\kappa_{\text{actor}}\in\{-0.99,-0.5,0,0.5,0.99\}$. The rationale for choosing more pessimistic (to semi-optimistic) $\kappa_{\text{critic}}$ values is to counteract the overestimation bias. Given that USAC stems from the framework of SAC, SAC was included in our figures for comparison.
Table 2: Best $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ parameters for USAC.
Environment $\kappa_{\text{critic}}$ $\kappa_{\text{actor}}$
Ant-v4 $-0.831559$ $-0.99$
HalfCheetah-v4 $-0.33$ $-0.50$
Hopper-v4 $-0.831559$ $0.50$
Humanoid-v4 $-0.831559$ $-0.831559^{\star}$
Walker2d-v4 $-0.831559$ $0.0$
By comparing the final returns (Figure 2) with the estimation error results (Figure 3) for the different grid combinations, we reaffirm previous findings [10, 7, 20]: i) overestimation (negative estimation error) should always be avoided as it reduces performance, and ii) underestimation (positive estimation error) does not necessarily harm performance. This indicates that a pessimistic approach should be adopted in critic training, whereas in actor training, either a pessimistic or optimistic approach can be suitable depending on the environment. The best combinations of $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$, with respect to final returns, are summarized in Table 2. Additionally, the area under the learning curve for the grid search parameters is illustrated in Figure 4 (Appendix C).
Figure 2: Grid search results for the final return of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes. The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
Figure 3: Grid search results for the estimation error (true - estimation) of USAC algorithm with Laplace distribution, averaged over three seeds, each with three evaluation episodes. The grid search explore various $\kappa_{\text{critic}}$ and $\kappa_{\text{actor}}$ values, using SAC as the baseline.
6 Discussion
We introduce Utility Soft Actor-Critic (USAC), a novel off-policy actor-critic algorithm. The key idea is to balance the trade-off between pessimism and optimism through incorporating a utility function that captures the uncertainty in the critic due to limited access to the environment. This framework is flexible and can be adapted to different distributions, with our experiments primarily focusing on the Laplace distribution. Unlike OAC [7] and TOP [20], which also explore optimism/pessimism in actor-critic training, USAC provides a policy improvement guarantee.
USAC allows independent control over the optimism and pessimism levels for the actor and critic. This decoupling enables a stable critic, necessary for accurate value function estimation, and an explorative actor, crucial for effective policy learning. We conducted a grid search experiment to explore these levels in training, revealing that being pessimistic in critic training while being either optimistic or pessimistic in actor training can be beneficial. However, it is also essential to consider the specific environment. Our work paves the way for further exploration in optimizing actor-critic algorithms with varying optimism and pessimism levels.
Limitations and future work
USAC’s primary limitation is its reliance on fixed optimism/pessimism parameters. An exciting direction for future research is to develop adaptive schemes where these parameters adjust dynamically throughout the learning process to better respond to the changing dynamics of the environment. There are several potential approaches to integrate automatic tuning. For example, these include: (i) adapting the automatic tuning scheme used for $\alpha$ in SAC [12, 13], (ii) employing bandit-based selection methods [20], or (iii) exploring continuous learning approaches using gradient descent [4].
Another promising direction for future work is to incorporate higher moments of the critic distributions. By considering not only the mean and variance but also skewness, kurtosis, and heavy-tailed distributions, we can better capture the nuances of uncertainty in the critic’s value estimates. This approach could lead to more robust and flexible policy learning, accommodating a wider variety of distributional characteristics. Incorporating these higher-order moments may provide a richer framework for addressing the biases and variances in the value function estimation.
Disclosure of funding
Appendix A Proofs
Proof of Proposition 1.
Suppose $Q(s,a)\sim\textrm{Laplace}(\mu_{\mathcal{Q}}(s,a),b_{\mathcal{Q}}(s,a))$. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected utility function of $\mathcal{Q}$ on $(s,a)$ can be expressed as
@@ Appendix A Proofs
At last, this expression can be further simplified using the definition of $g(\kappa)$ in 12.
Appendix B USAC under gaussianity

### v2

@@ 4 A utility-theoretic framework to make the action-value biases steerable
where $U^{\mathcal{Q}}_{\lambda}(s,a)=\mathbb{E}_{Q\sim\mathcal{Q}}[Q(s,a)]$ when $\lambda=0$. This utility function is commonly used for measuring risk in finance, economics, and decision-making under uncertainty [11, 37]. It also shares similarities with the utility functions employed in risk-sensitive RL [16, 29, 25] and distributional RL [27, 3, 4]. It is noteworthy that we apply this utility function to the $\mathcal{Q}$ distribution rather than the return distribution, as is typically done in risk-sensitive and distributional RL. The utility function is also referred to as expected utility, exponential utility, exponential risk measure, generalized mean, or entropic risk measure according to the context [10, 11].
@@
We use ˜8 as a measure of pessimism or optimism of the $\mathcal{Q}$ distribution: a positive value of $\lambda$ implies an optimistic view, while a negative value indicates a pessimistic view. This interpretation is clear from a second-order Taylor expansion applied to the utility function:
$U^{\mathcal{Q}}_{\lambda}=\mathbb{E}_{Q\sim\mathcal{Q}}[Q]+\frac{\lambda}{2}\mathbb{V}_{Q\sim\mathcal{Q}}[Q]+\mathcal{O}(\lambda^{2}).$
Consequently, the utility function $U^{\mathcal{Q}}_{\lambda}$ spans the entire spectrum of the $\mathcal{Q}$ distribution by varying $\lambda$.
4.1 Policy evaluation, improvement and guarantees
For any given policy $\pi\in\Pi$, the utility $U^{\mathcal{Q}}_{\lambda}\in\mathcal{B}(\mathcal{S}\times\mathcal{A})$ will reduce to $Q^{\pi}$ for any $|\lambda|<\infty$, when $\mathcal{Q}$ is a Dirac delta distribution centered at $Q^{\pi}$. This implies that the deviation of $U^{\mathcal{Q}}_{\lambda}$ from $Q^{\pi}$ is due to the uncertainty in the learned $Q$-function. In fact, we observe that the Bellman operator should also bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed point $Q^{\pi}$ [5]
Lemma 1.
For any expected utility of $\mathcal{Q}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$, we have
@@ 4.1 Policy evaluation, improvement and guarantees
$\lVert T^{\pi}U^{\mathcal{Q}}_{\lambda}-U^{\mathcal{Q}}_{\lambda}\rVert^{2}_{\rho^{\pi}}\leq\mathbb{E}_{(s,a)\sim\rho^{\pi}}\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[\ell(U^{\mathcal{Q}}_{\lambda};\xi)\right],$
where
$\ell(U^{\mathcal{Q}}_{\lambda};\xi)=(r(s,a)+\gamma(U^{\mathcal{Q}}_{\lambda}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime})))-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}$
is the squared loss of $U^{\mathcal{Q}}_{\lambda}$ on an observation $\xi=(s,a,s^{\prime},a^{\prime})$.
Proof.
For any $\mathcal{Q}\in\mathcal{P}(\mathcal{S}\times\mathcal{A})$, we have
$\lVert T^{\pi}U^{\mathcal{Q}}_{\lambda}-U^{\mathcal{Q}}_{\lambda}\rVert_{\rho^{\pi}}^{2}=\mathbb{E}_{(s,a)\sim\rho^{\pi}}[(T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}]$
$=\mathbb{E}_{(s,a)\sim\rho^{\pi}}[(r(s,a)+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[U^{\mathcal{Q}}_{\lambda}(s^{\prime},a^{\prime})$-
$\quad\quad\quad\quad\quad\quad\alpha\log(\pi(a^{\prime}|s^{\prime}))]-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}]$
$\leq\mathbb{E}_{(s,a)\sim\rho^{\pi}}\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}[(r(s,a)+\gamma(U^{\mathcal{Q}}_{\lambda}(s^{\prime},a^{\prime})$-
$\quad\quad\quad\quad\quad\quad\alpha\log(\pi(a^{\prime}|s^{\prime})))-U^{\mathcal{Q}}_{\lambda}(s,a))^{2}],$
where the inequality comes from the Jensen’s inequality. ∎
Thus, as a consequence of Lemma˜1, the Bellman error defined by
$T^{\pi}U^{\mathcal{Q}}_{\lambda}(s,a)-U^{\mathcal{Q}}_{\lambda}(s,a)=r(s,a)$
$\qquad+\gamma\mathbb{E}_{(s^{\prime},a^{\prime})\sim p^{\pi}(\cdot|s,a)}\left[U_{\lambda}^{\mathcal{Q}}(s^{\prime},a^{\prime})-\alpha\log(\pi(a^{\prime}|s^{\prime}))\right]-U_{\lambda}^{\mathcal{Q}}(s,a)$
@@
should diminish. Nevertheless, it fosters exploration during learning by maintaining a distribution over the $Q$-functions. On the other hand, given a $\mathcal{Q}$ distribution associated with the current policy $\pi$, denoted as $\mathcal{Q}^{\pi}$, we update the policy, similar to ˜4, through the following maximization problem:
@@
4.2 The Utility Soft Actor Critic (USAC)
We illustrate a simple implementation of our utility-theoretic formulation into a soft actor-critic pipeline and refer to the resulting algorithm as the Utility Soft Actor Critic (USAC). We perform online and off-policy learning as usual, maintaining a replay buffer of observed environment interactions $(s,a,r,s^{\prime})\sim\mathcal{D}$. As the operator $T^{\pi}$ should bring the utility $U^{\mathcal{Q}}_{\lambda}$ to its fixed-point $Q^{\pi}$, the USAC framework learns the critic $Q$ (an estimate of $Q^{\pi}$) as in ˜2 but with the critic’s target value defined as
$y(s,a,r,s^{\prime})=r+\gamma[U_{\lambda_{\text{critic}}}^{\bar{\mathcal{Q}}}(s^{\prime},a^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})],$ (9)
where $a^{\prime}\sim\pi(\cdot|s^{\prime})$. Note that we use a target critic distribution $\bar{\mathcal{Q}}$ with some utility parameter $\lambda_{\text{critic}}$ chosen specifically for the critic. Similar to ˜4, the actor learns its policy by
@@ 4.2 The Utility Soft Actor Critic (USAC)
Note that Moskovitz et al. [22] implements a single shared optimism/pessimism parameter for the actor and the critic. In contrast, our framework naturally decouples the optimism and pessimism levels of the actor and critic into two separate scalars, $\lambda_{\text{critic}}$ and $\lambda_{\text{actor}}$, respectively. This decoupling allows both investigating the relationship between the optimism and pessimism demands of the actor and the critic and tuning them separately, thereby benefiting from their complementary roles in online training. Later in the experiments presented in Section˜5, we will explore various combinations of optimism and pessimism levels.
Algorithm 1 USAC with Laplace distribution (Section˜4.3)
Averaging parameter $\tau\in(0,1)$, learning rates $\eta_{\theta},\eta_{\phi}>0$, mini-batch size $n\in\mathbb{N}$, pessimism/optimism coefficients $\kappa_{\text{actor}},\kappa_{\text{critic}}\in(-1,1)$, entropy coefficient $\alpha$,
1:Polyak coefficient $\tau$
[…]
4.3 Laplace distributed critics
The utility formulation in ˜8 enables us to compute the utility for any distribution with a moment-generating function. We will next present an implementation of a specific case of our USAC algorithm where $\mathcal{Q}$ follows a Laplace distribution, which has the desirable property of expressing the utility parameter within the [1,+1][-1,+1]. The pseudo-code of the resulting USAC algorithm is provided in Algorithm˜1. Here, we denote the utility parameter associated with this distribution as $\kappa$, reflecting its characteristics within the Laplace context.
@@ 4.3 Laplace distributed critics
[…]
Proof.
Suppose $Q(s,a)\sim\textrm{Laplace}(\mu_{\mathcal{Q}}(s,a),b_{\mathcal{Q}}(s,a))$. Then, for any $(s,a)\in\mathcal{S}\times\mathcal{A}$, the expected utility function of $\mathcal{Q}$ on $(s,a)$ is
@@
At last, this expression can be further simplified using the definition of $g(\kappa)$ in ˜12. ∎
In Proposition˜2, we observe that by varying $\kappa$ in the small interval (1,1)(-1,1), $U^{\mathcal{Q}}_{\kappa}$ spans the entire spectrum of the $\mathcal{Q}$ distribution. Specifically, as $\lim_{\kappa\rightarrow-1}g(\kappa)\rightarrow$-$\infty$, $\lim_{\kappa\rightarrow 0}g(\kappa)\rightarrow 0$, and $\lim_{\kappa\rightarrow 1}g(\kappa)\rightarrow\infty$. Alternatively, one could also presume Gaussian-distributed $Q$-values.